OpenStudy (anonymous):

Complete the equation of the graphed linear function in point-slope form. y – (–2) = ? (x – ?)

OpenStudy (solomonzelman):

what information are you given so that you can fill in blanks

OpenStudy (anonymous):

I'm attaching the graph now.

OpenStudy (solomonzelman):

But you know that this is a point slope form. \(\large\color{black}{ (y-y_1)=m(x-x_1) }\) where the point is (x1,y1) and m is the slope.

OpenStudy (anonymous):

I have to fill in the blank, which is where the question marks are.

OpenStudy (solomonzelman):

Lets visualize the slope. lets look at a change in x. when x is 1, the y is -2. when x is 2, the y is 2. So the slope is 4.

OpenStudy (solomonzelman):

Visualising.

OpenStudy (solomonzelman):

So, when slope is 4, and the point is (1,-2) fill in blanks. \(\large\color{black}{ (y-(-2)~~)=4(x-(1)~~) }\)

OpenStudy (solomonzelman):

Just another a-(-b) = a+b to finish the problem.